The radius of the moon is one-fourth, and its mass is one eighty-first that of the earth. If the
acceleration due to gravity on the surface of the earth is 9.8 m/s2, what is its value on the
moon’s surface.
1
Expert's answer
2019-04-26T11:44:28-0400

The acceleration due to gravity


"g=G\frac{M}{R^2}=G\frac{M_{\bigoplus}/81}{\left(R_{\bigoplus}/4\right)^2}=\frac{16}{81}g_{\bigoplus}"

"=\frac{16}{81}\times 9.8=1.9\:\rm{m/s^2}"